1/3(c+1)=1/6(3c-5)
Answer by checkley77(12844)

1/3(c+1)=1/6(3c-5)
(C+1)/3=(3C-5)/6 CROSS MULTIPLY
3(3C-5)=6(C+1)
9C-15=6C+6
9C-6C=6+15
3C=21
C=21/3
C=7 ANSWER.
PROOF:
(7+1)/3=(3*7-5)/6
8/3=(21-5)/6
8/3=16/6
48=48